# Can We Detect Failures Without Failure Data? Uncertainty-Aware Runtime Failure Detection for Imitation Learning Policies

### v1

@@ II Related Work
OOD Detection. The task of detecting robot failures can be viewed as anomaly detection, which falls under the broader framework of OOD detection [61]. Ensemble methods [31], which combine predictions from multiple models to improve robustness and estimate uncertainty, have long been regarded as the de facto approach for addressing this problem. However, they are computationally expensive as they require training and running inference on multiple models. Another popular approach frames OOD detection as a classification problem [34]. This formulation learns the decision boundary between ID and OOD data by training a binary classifier [54, 16], but requires OOD data during training. In contrast, density-based approaches [35, 17, 60], one-class discriminators based on random networks [13, 22], and control-theoretic methods [7] aim to model information from ID data without relying on OOD data during training. Density-based methods attempt to capture the distribution of ID data, yet they can be challenging to optimize. One-class discriminators have shown superior performance over deep ensembles in practice but can be sensitive to the design of the discriminator model. Control-theoretic approaches use contrastive energy-based models; however, they often require a representation of the system’s dynamics. Furthermore, evidential deep learning methods [8, 5, 26] learn parameters for second-order distributions (e.g., Dirichlet) to approximate epistemic uncertainty (due to limited model knowledge or OOD inputs) from aleatoric uncertainty (due to inherent randomness in the data). Lastly, distance-based approaches [3, 52] identify OOD samples by computing their distance to ID samples in the input or latent space, avoiding the need for training but exhibiting limited performance compared to other approaches. We consider many of the listed model variants as score candidates in FAIL-Detect.
@@
Failure Detection in Robotics. Detecting failures in robotic systems is important for ensuring safety and reliability, as failures can lead to undesirable behaviors in human environments [39, 42, 34]. Various approaches have been proposed, such as building fast anomaly classifiers based on LLM embeddings [49] and using the reconstruction error from variational autoencoders (VAE) to detect anomalies in behavior cloning (BC) policies for mobile manipulation [57]. Separately, Ren et al. [43] construct uncertainty sets from conformal prediction for actions generated by an LLM-based planner, prompting human intervention when the set is ambiguous. These works do not consider failure detection in the setting of generative imitation learning policies. On the other hand, Gokmen et al. [19] learn a state value function that is trained jointly with a BC policy and can be used to predict failures. Liu et al. [33] propose an LSTM-based failure classifier for a BC-RNN policy using latent embeddings from a conditional VAE. Given a Transformer-based policy and a world model to predict future latent embeddings, Liu et al. [34] train a failure detection classifier on the embeddings. To handle previously unseen states, they also propose a SOTA OOD detection method, which we adapt as a baseline to our approach (see PCA-kmeans in Section V). However, unlike FAIL-Detect, these methods require collecting failed trajectories a priori to detect failures. For diffusion-based policies, Sun et al. [51] reduce model uncertainty by producing prediction intervals for rewards of predicted trajectories. He et al. [22] propose using random network distillation (RND) to detect OOD trajectories and select reliable ones. These works do not directly consider runtime failure detection. Our two-stage solution for this problem combines the advantages of both approaches. The closest SOTA method to FAIL-Detect by Agia et al. [1] introduces a statistical temporal action consistency (STAC) measure in conjunction with vision-language models (VLMs) to detect failures within rollouts at runtime. STAC does not require failure data, consists of a score computed post-hoc from a batch of predicted actions and a constant-time CP threshold to flag failures, and is evaluated in the context of DP. We demonstrate improved empirical performance over STAC by integrating learned failure detection scores with a time-varying CP band.
@@
TABLE I: Overview of score methods evaluated in this work. The input was selected either based on the structure and requirements of each method or, when multiple input combinations were possible, based on empirical performance.
@@ IV-A Design of Scalar Scores
(a) Learned data density: we fit a normalizing flow-based density estimator to the observations, where data far from the distribution of successful trajectory observations may indicate failure. The approach we term lopO [60] fits a continuous normalizing flow (CNF) $f_{\theta}$ to the set of observations $\{O_{t}\}_{t\geq 0}$. A low $\log p(O_{t^{\prime}})$ for a new observation $O_{t^{\prime}}$ implies it is unlikely, indicating possible failure. Note the computation of $\log p(O_{t^{\prime}})$ requires integration of the divergence of $f_{\theta}$ over the ODE trajectory, which is difficult to estimate in high dimensions. Additionally, we introduce our novel logpZO approach, which leverages the same CNF $f_{\theta}$ to evaluate the likelihood of a noise estimate $Z_{O_{t}}$ (conditioned on an observation $O_{t}$). Using the forward ODE process, we compute $Z_{O_{t}}$ by integrating $f_{\theta}$ over the unit interval [0,1], starting from $O_{t}$ as the ODE initial condition. When $O_{t}$ is ID, $Z_{O_{t}}$ is approximately Gaussian, leading to $p(Z_{O_{t}})=C\exp(-0.5|Z_{O_{t}}|^{2})$. Thus, a high value of $|Z_{O_{t}}|_{2}^{2}$ corresponds to a low likelihood $p(Z_{O_{t}})$ in the noise space. The key distinction between lopO and logpZO lies in their domains: the former assesses likelihood in the original observation space, while the latter does so in the latent noise space. We expect the latter to be better because its computation does not require the divergence of $f_{\theta}$ integrated over $[0,1]$, a hard-to-estimate quantity in high dimensions.
@@
Figure 4: Quantitative failure detection results for simulation tasks on FM policy ( best, second, third); results with TPR and TNR are in Fig. 9 and results on DP are in Fig. 10. For accuracy and weighted accuracy, higher is better and for detection time, lower is better. The CP band for each task is calibrated with successful rollouts under ID initial conditions only (i.e., the same band is used for ID and OOD test cases). We group together post-hoc (STAC, PCA-kmeans, SPARC), density-based (logpO, logpZO), second-order (DER, NatPN), and one-class (CFM, RND) methods and show barplots with standard errors. The dashed line in the Detection Time plots represents the average successful trajectory time in that setting with standard error. Overall, learned methods outperform post-hoc ones in failure detection. In terms of combined accuracy (accuracy and weighted accuracy), both logpZO and RND perform the best by reaching top-3 performance in 12/16 cases. Moreover, logpZO reaches the best overall performance in 10/16 cases, while RND does so in 4/16 cases. In comparison, the baselines STAC and PCA-kmeans only reach top-3 performance in 6/16 and 3/16 cases, respectively. Note that STAC reaches the best overall performance in 3/16 cases, while PCA-kmeans never does. The learned methods also achieve the fastest detection time with one of the learned methods always getting the best overall detection time in all eight cases. In terms of best top-3 performance, logpZO achieves it in 7/8 cases. RND detects failures before the average success end time, but is slower to detect than logpZO with top-3 performance in only 3/8 cases. The PCA-kmeans baseline ranks in the top-3 detection time in 5/8 cases. In contrast, STAC is the slowest in nearly all cases, detecting failures only after the average success trajectory time, rendering the detection not practical.
@@
Figure 5: Quantitative results for the robot hardware experiments across two tasks with policies trained using FM and DP. We consider two different ways to compute the CP band: “setting-dependent” using successful trajectories from each OOD/ID environment and “ID-only” using only the trajectories from the ID environment. For accuracy and weighted accuracy, higher is better and for detection time, lower is better. Additional metrics are reported in Fig. 11. The figure layout is the same as Fig. 4 ( best, second, third), and ‘NaN’ detection time indicates that no test rollout was detected as failed. Once again the learned approaches outperform the post-hoc methods. Note we do not present STAC here as it was slow to run on hardware in real-time. At the small sample size regime, logpZO remains robust in combined accuracy, achieving top-3 performance in 10/12 cases and overall best performance in the highest number of cases (5/12). In contrast, the PCA-kmeans baseline reaches top-3 performance in 9/12 cases and is overall best in only 2/12 cases. The learned CFM score ties with PCA-kmeans in performance and is best in 3/12 cases. RND follows with 8/12 cases in top-3 performance. In detection time, the learned methods attained fastest detection time in 5/6 cases. Despite having lower accuracy and usually not achieving the fastest time, PCA-kmeans is robust in speed as it attains top-3 performance in 5/6 cases. logpZO still remains practical with detection times well below the average success trajectory completion time.
@@ IV-B Sequential Threshold Design with Conformal Prediction
For sequential failure detection, we build $C_{\alpha}$ as a one-sided time-varying CP band. The band is one-sided as we are only concerned with high values of the scalar score $D_{M}(A_{t},O_{t};\theta)$, which indicate the trajectory is OOD (i.e., a failure). Given $N$ successful rollouts as the calibration data, we obtain scalar scores $\mathcal{D}_{cal}=\{D_{M}(A_{t}^{i},O_{t}^{i};\theta):i=1,\ldots,N\text{ and }t=1,H^{\prime},\ldots,T\}$. The CP band is a set of intervals $C_{\alpha}=\{[\text{lower}_{t},\text{upper}_{t}]:t=1,H^{\prime},\ldots,T\}$, where $\text{lower}_{t}~{}\equiv~{}\min(\mathcal{D}_{cal})$ since the band is one-sided. To obtain the upper bound, we follow [15], computing the time-varying mean $\mu_{t}$ and band width $h_{t}$, so that $\text{upper}_{t}=\mu_{t}+h_{t}$. Further details of upper bound construction are in Section -A. Theoretically, for a new successful rollout $\tau_{T}=(O_{0},A_{0},\ldots,O_{T},A_{T})$, with probability at least $1-\alpha$, the score $D_{M}(A_{t},O_{t};\theta)\in[\text{lower}_{t},\text{upper}_{t}]$ for all $t=1,H^{\prime},\ldots,T$. By defining the threshold $\eta_{t}=\text{upper}_{t}$ and setting failures to one, the decision rule $\mathbbm{1}(D_{M}(A_{t},O_{t};\theta)>\eta_{t})$ controls the false positive rate (successes marked as failures) at level $\alpha$.
@@ V Experiments
We test our two-stage failure detection framework in both simulation and on robot hardware. Our experiments span multiple environments, each presenting unique challenges in terms of types of tasks and distribution shifts. We empirically investigate an extensive set of both learned and post-hoc scalar scores (see Table I) within our FAIL-Detect framework (see results in Section VI). We refer to Section -B for more details on policy training, the CP band calibration procedure, and the learned scalar score architectures.
@@
To quantify failure detection performance, we denote failed rollouts as one and successful rollouts as zero. We then adopt the following standard metrics: (1) true positive rate (TPR), (2) true negative rate (TNR), (3) accuracy = (TPR + TNR) / 2, (4) weighted accuracy = $\beta\cdot$TPR + $(1-\beta)\cdot$ TNR for $\beta=\frac{\#\text{Successful rollouts}}{\#\text{Rollouts}}$, and (5) detection time = $\mathbb{E}_{(A_{t},O_{t})\sim\text{test rollouts}}[\arg\min_{t=1,H^{\prime},\ldots,T}\mathbbm{1}(D_{M}(A_{t},O_{t};\theta)>\eta_{t})]$, which computes the average failure detection time from the start of the rollout. The accuracy metric equally represents classes in an imbalanced dataset (e.g., few successful rollouts in an OOD setting). Weighted accuracy represents how well a method matches the true success / failure distribution. Due to the high human time cost of performing real-robot rollouts, we evaluate FAIL-Detect and the baselines on significantly fewer rollouts in the robot hardware tasks (i.e., 50 rollouts) compared to the simulation tasks (i.e., 2000 rollouts).
@@ VI-A How performant is failure detection without failure data?
A key question we consider is whether failure detection is possible and performant without enumerating all possible failure scenarios, which is practically infeasible. We conduct extensive experiments across simulation and robot hardware tasks to answer this question. We evaluate accuracy, weighted accuracy, and detection time to assess whether failures can be identified reliably and quickly.
FAIL-Detect achieves high accuracy with fast detection. Our two-stage framework demonstrates strong performance across both accuracy metrics and detection speed. For example, the average best accuracy across FAIL-Detect’s score candidates is $\sim 75\%$ in simulation (Fig. 4) and $\sim 74\%$ on the robot hardware tasks (Fig. 5). This performance shows the capacity of failure-free failure detection methods to robustly identify failures across many scenarios. Notably, FAIL-Detect maintains viable detection time across various score designs, with average best detection time faster than successful trajectory completion.
@@
Learned scores outperform post-hoc scores. Looking at performance across simulation and robot hardware tasks, we find that learned scalar scores hold an advantage over post-hoc scores in failure detection. In simulation (Fig. 4), logpZO and RND perform best across all accuracy metrics, achieving top-3 performance in 12/16 cases. logpZO further reaches the best overall accuracy in 10/16 cases, while RND does so in 4/16 cases. Specifically, logpZO outperforms in weighted accuracy, better representing the true failure/success distribution. In comparison, RND better balances high performance in both the success and failure categories individually, attaining high accuracy. Among the considered post-hoc scores, SPARC as part of FAIL-Detect outperforms the STAC and PCA-kmeans baselines with top-3 performance in 7/16 cases. STAC comes in second in the post-hoc category at top-3 in 6/16 cases. Note that STAC achieves the best overall accuracy in 3/16 yet PCA-kmeans never does. Overall, there is a large performance gap between the learned and post-hoc methods. We did notice that post-hoc methods perform better in the OOD cases than in ID scenarios. We hypothesize this may be due to a clearer distinction between successful ID trajectories and failed OOD trajectories, for example, for SPARC if the OOD trajectory exhibits significant jitter.
In detection time, logpZO is most efficient, ranking among the three fastest methods in 7/8 cases and as the fastest method in 3/8 cases. Despite RND’s high accuracy performance, it is a little slower to detect, but remains faster than the average time to success and is the fastest in 1/8 cases. Notably, STAC’s detection time consistently exceeds practical limits, surpassing the average success trajectory time. In contrast, PCA-kmeans has a top-3 detection time in 5/8 cases, but does not come with high failure detection accuracy.
@@
For the robot hardware experiments (Fig. 5), with a much lower rollout data regime for calibration than in simulation and a wider diversity of behaviors and observations, logpZO remains robust as the best method, reaching top-3 highest accuracy and weighted accuracy in 10/12 scenarios. RND and PCA-kmeans perform comparably with 8/12 and 9/12 top-3 rankings, respectively. Additionally, most methods, including logpZO and RND, maintain practical detection times well below the average successful trajectory time.
Overall, across all experiments, we find our novel learned logpZO score within the FAIL-Detect framework to be most consistent in performance. Post-hoc methods are often at the extremes of performance (either doing well or poorly) depending on the particular setting.
Qualitative score trends. Visualization of the detection scores (Figs. 6 and 12) confirms that learned methods are more discriminative with better score separation between successful and failed trajectories compared to post-hoc approaches. STAC also suffers from a single calibration threshold that is time invariant. In Section -C, we present comprehensive ablation studies examining: (1) STAC-adaptive, our modification replacing STAC’s constant CP band with a time-varying CP band (Table IV), (2) performance sensitivity to CP significance level $\alpha$ (Table V), and (3) performance across different model architectures (Fig. 13).

### v3

@@ II Related Work
OOD Detection. The task of detecting robot failures can be viewed as anomaly detection, which falls under the broader framework of OOD detection [61]. Ensemble methods [31], which combine predictions from multiple models to improve robustness and estimate uncertainty, have long been regarded as the de facto approach for addressing this problem. However, they are computationally expensive as they require training and running inference on multiple models. Another popular approach frames OOD detection as a classification problem [34]. This formulation learns the decision boundary between ID and OOD data by training a binary classifier [53, 15], but requires OOD data during training. In contrast, density-based approaches [35, 16, 60], one-class discriminators based on random networks [13, 21], and control-theoretic methods [7] aim to model information from ID data without relying on OOD data during training. Density-based methods attempt to capture the distribution of ID data, yet they can be challenging to optimize. One-class discriminators have shown superior performance over deep ensembles in practice but can be sensitive to the design of the discriminator model. Control-theoretic approaches use contrastive energy-based models; however, they often require a representation of the system’s dynamics. Furthermore, evidential deep learning methods [8, 5, 25] learn parameters for second-order distributions (e.g., Dirichlet) to approximate epistemic uncertainty (due to limited model knowledge or OOD inputs) from aleatoric uncertainty (due to inherent randomness in the data). Lastly, distance-based approaches [3, 51, 27] identify OOD samples by computing their distance to ID samples in the input or latent space, avoiding the need for training but exhibiting limited performance compared to other approaches. We consider many of the listed model variants as score candidates in FAIL-Detect.
@@
Failure Detection in Robotics. Detecting failures in robotic systems is important for ensuring safety and reliability, as failures can lead to undesirable behaviors in human environments [39, 42, 34]. Various approaches have been proposed, such as building fast anomaly classifiers based on LLM embeddings [48] and using the reconstruction error from variational autoencoders (VAE) to detect anomalies in behavior cloning (BC) policies for mobile manipulation [57]. Separately, Ren et al. [43] construct uncertainty sets from conformal prediction for actions generated by an LLM-based planner, prompting human intervention when the set is ambiguous. These works do not consider failure detection in the setting of generative imitation learning policies. On the other hand, Gokmen et al. [18] learn a state value function that is trained jointly with a BC policy and can be used to predict failures. Liu et al. [33] propose an LSTM-based failure classifier for a BC-RNN policy using latent embeddings from a conditional VAE. Given a Transformer-based policy and a world model to predict future latent embeddings, Liu et al. [34] train a failure detection classifier on the embeddings. To handle previously unseen states, they also propose a SOTA OOD detection method, which we adapt as a baseline to our approach (see PCA-kmeans in Section V). However, unlike FAIL-Detect, these methods require collecting failed trajectories a priori to detect failures. Meanwhile, Wang et al. [55] uses self-reset to collect additional failure data and train a classifier to identify failure modes. In their on-robot experiments, approximately 2000 trajectories (roughly 2 hours) had to be collected using self-reset, making scalability challenging. For diffusion-based policies, Sun et al. [50] reduce model uncertainty by producing prediction intervals for rewards of predicted trajectories. He et al. [21] propose using random network distillation (RND) to detect OOD trajectories and select reliable ones. These works do not directly consider runtime failure detection. Our two-stage solution for this problem combines the advantages of both approaches. The closest SOTA method to FAIL-Detect by Agia et al. [1] introduces a statistical temporal action consistency (STAC) measure in conjunction with vision-language models (VLMs) to detect failures within rollouts at runtime. STAC does not require failure data, consists of a score computed post-hoc from a batch of predicted actions and a constant-time CP threshold to flag failures, and is evaluated in the context of DP. We demonstrate improved empirical performance over STAC by integrating learned failure detection scores with a time-varying CP band.
@@
TABLE I: Overview of score methods evaluated in this work. The input was selected either based on the structure and requirements of each method or, when multiple input combinations were possible, based on empirical performance. All methods except STAC (which proposes a different calibration method; see Section V) use time-varying CP bands described in Section IV-B.
@@ IV-A Design of Scalar Scores
(a) Learned data density: we fit a normalizing flow-based density estimator to the observations, where data far from the distribution of successful trajectory observations may indicate failure. The approach we term lopO [60] fits a continuous normalizing flow (CNF) $f_{\theta}$ to the set of observations $\{O_{t}\}_{t\geq 0}$. A low $\log p(O_{t^{\prime}})$ for a new observation $O_{t^{\prime}}$ implies it is unlikely, indicating possible failure. Note the computation of $\log p(O_{t^{\prime}})$ requires integration of the divergence of $f_{\theta}$ over the ODE trajectory, which is difficult to estimate in high dimensions. Additionally, we introduce our novel logpZO approach, which leverages the same CNF $f_{\theta}$ to evaluate the likelihood of a noise estimate $Z_{O_{t}}$ (conditioned on an observation $O_{t}$). Using the forward ODE process, we compute $Z_{O_{t}}$ by integrating $f_{\theta}$ over the unit interval [0,1], starting from $O_{t}$ as the ODE initial condition. When $O_{t}$ is ID, $Z_{O_{t}}$ is approximately Gaussian, leading to $p(Z_{O_{t}})=C\exp(-0.5|Z_{O_{t}}|^{2})$. Thus, a high value of $|Z_{O_{t}}|_{2}^{2}$ corresponds to a low likelihood $p(Z_{O_{t}})$ in the noise space. Further details on logpZO are described in Section -A. The key distinction between lopO and logpZO lies in their domains: the former assesses likelihood in the original observation space, while the latter does so in the latent noise space. We expect the latter to be better because its computation does not require the divergence of $f_{\theta}$ integrated over $[0,1]$, a hard-to-estimate quantity in high dimensions.
@@
Figure 4: Quantitative failure detection results for simulation tasks on FM policy ( best, second, third); results with TPR and TNR are in Fig. 11 and results on DP are in Fig. 12. For balanced accuracy and weighted accuracy, higher is better and for detection time, lower is better. The CP band for each task is calibrated with successful rollouts under ID initial conditions only (i.e., the same band is used for ID and OOD test cases). We group together post-hoc (STAC, PCA-kmeans, SPARC), density-based (logpO, logpZO), second-order (DER, NatPN), and one-class (CFM, RND) methods and show barplots with standard errors. The dashed line in the Detection Time plots represents the average successful trajectory time in that setting with standard error. Overall, learned methods outperform post-hoc ones in failure detection. In terms of combined accuracy (balanced accuracy and weighted accuracy), logpZO and RND are the best two methods, reaching top-1 performance in 10/16 and 5/16 cases, respectively. Moreover, logpZO reaches top-3 performance in 14/16 cases, while RND does so in 9/16 cases. In comparison, the baselines STAC and PCA-kmeans reach top-1 performance in 3/16 and 0/16 cases, respectively. Note that STAC reaches top-3 performance in 8/16 cases, while PCA-kmeans does so in 3/16 cases. The learned methods also achieve the fastest detection time, with one of the learned methods always getting the best overall detection time in all but one case. In terms of best top-1 performance, logpZO is the fastest method in 3/8 cases, RND in 0/8 cases, and the PCA-kmeans baseline does so in 1/8 cases. In contrast, STAC is the slowest in nearly all cases, detecting failures only after the average success trajectory time, rendering the detection not practical.
@@
Figure 5: Quantitative results for the robot hardware experiments across two tasks with policies trained using FM and DP. We consider two different ways to compute the CP band: “setting-dependent” using successful trajectories from each OOD/ID environment and “ID-only” using only the trajectories from the ID environment. For balanced accuracy and weighted accuracy, higher is better and for detection time, lower is better. Additional metrics are reported in Fig. 13 and Fig. 14. The figure layout is the same as Fig. 4 ( best, second, third), and ‘NaN’ detection time indicates that no test rollout was detected as failed. Once again the learned approaches outperform the post-hoc methods. Note we do not present STAC here as it was slow to run on hardware in real-time. In the small sample size regime, logpZO remains robust in combined accuracy, achieving top-1 performance in the highest number of cases (8/12) and top-3 performance in 11/12 cases. RND underperforms by never reaching top-1 performance, yet it always achieves top-3 performance. In contrast, the PCA-kmeans baseline reaches top-1 performance in 4/12 cases and top-3 performance in 10/12 cases. In detection time, the post-hoc SPARC method is the fastest in 4/6 cases, yet it never achieves top-1 performance. PCA-kmeans is robust in speed as it attains top-3 performance in 4/6 cases. logpZO still remains practical with detection times well below the average success trajectory completion time.
@@ IV-B Sequential Threshold Design with Conformal Prediction
For sequential failure detection, we build $C_{\alpha}$ as a one-sided time-varying CP band. The band is one-sided as we are only concerned with high values of the scalar score $D_{M}(A_{t},O_{t};\theta)$, which indicate the trajectory is OOD (i.e., a failure). Given $N$ successful rollouts as the calibration data, we obtain scalar scores $\mathcal{D}_{cal}=\{D_{M}(A_{t}^{i},O_{t}^{i};\theta):i=1,\ldots,N\text{ and }t=1,H^{\prime},\ldots,T\}$. The CP band is a set of intervals $C_{\alpha}=\{[\text{lower}_{t},\text{upper}_{t}]:t=1,H^{\prime},\ldots,T\}$, where $\text{lower}_{t}~{}\equiv~{}\min(\mathcal{D}_{cal})$ since the band is one-sided. To obtain the upper bound, we follow [14], computing the time-varying mean $\mu_{t}$ and band width $h_{t}$, so that $\text{upper}_{t}=\mu_{t}+h_{t}$. Further details of upper bound construction are in Section -B. Theoretically, for a new successful rollout $\tau_{T}=(O_{0},A_{0},\ldots,O_{T},A_{T})$, with probability at least $1-\alpha$, the score $D_{M}(A_{t},O_{t};\theta)\in[\text{lower}_{t},\text{upper}_{t}]$ for all $t=1,H^{\prime},\ldots,T$. By defining the threshold $\eta_{t}=\text{upper}_{t}$ and setting failures to one, the decision rule $\mathbbm{1}(D_{M}(A_{t},O_{t};\theta)>\eta_{t})$ controls the false positive rate (successes marked as failures) at level $\alpha$.
@@ V Experiments
We test our two-stage failure detection framework in both simulation and on robot hardware. Our experiments span multiple environments, each presenting unique challenges in terms of types of tasks and distribution shifts. We empirically investigate an extensive set of both learned and post-hoc scalar scores (see Table I) within our FAIL-Detect framework (see results in Section VI). We refer to Section -C for more details on policy training, the CP band calibration procedure, and the learned scalar score architectures.
@@
To quantify failure detection performance, we denote failed rollouts as one and successful rollouts as zero. We then adopt the following standard metrics: (1) true positive rate (TPR), (2) true negative rate (TNR), (3) balanced accuracy = (TPR + TNR) / 2, (4) weighted accuracy = $\beta\cdot$TPR + $(1-\beta)\cdot$ TNR for $\beta=\frac{\#\text{Successful rollouts}}{\#\text{Rollouts}}$, and (5) detection time = $\mathbb{E}_{(A_{t},O_{t})\sim\text{test rollouts}}[\arg\min_{t=1,H^{\prime},\ldots,T}\mathbbm{1}(D_{M}(A_{t},O_{t};\theta)>\eta_{t})]$, which computes the average failure detection time from the start of the rollout. The balanced accuracy metric equally represents classes in an imbalanced dataset (e.g., few successful rollouts in an OOD setting). Weighted accuracy represents how well a method matches the true success / failure distribution. Due to the high human time cost of performing real-robot rollouts, we evaluate FAIL-Detect and the baselines on significantly fewer rollouts in the robot hardware tasks (i.e., 50 rollouts) compared to the simulation tasks (i.e., 2000 rollouts).
@@ VI-A How performant is failure detection without failure data?
A key question we consider is whether failure detection is possible and performant without enumerating all possible failure scenarios, which is practically infeasible. We conduct extensive experiments across simulation and robot hardware tasks to answer this question. We evaluate balanced accuracy, weighted accuracy, and detection time to assess whether failures can be identified reliably and quickly.
[…]
@@
Learned scores outperform post-hoc scores. Looking at performance across simulation and robot hardware tasks, we find that learned scalar scores hold an advantage over post-hoc scores in failure detection. In simulation (Fig. 4), logpZO and RND are the best two methods, achieving top-1 performance in 10/16 and 5/16 cases, respectively. STAC is the best in the post-hoc category for top-1 accuracy in 3/16 cases, yet PCA-kmeans is never the best. Overall, there is a large performance gap between the learned and post-hoc methods, especially in terms of the best overall accuracy. We did notice that post-hoc methods perform better in the OOD cases than in ID scenarios. We hypothesize this may be due to a clearer distinction between successful ID trajectories and failed OOD trajectories, for example, for SPARC if the OOD trajectory exhibits significant jitter.
In terms of detection time, logpZO is the most efficient, achieving the fastest time in 3/8 cases, while PCA-kmeans does so in only 1/8 cases. Notably, STAC’s detection time consistently exceeds practical limits, surpassing the average success trajectory time.
@@
For the robot hardware experiments (Fig. 5), with a much lower rollout data regime for calibration than in simulation (see Table III) and a wider diversity of behaviors and observations, logpZO remains robust as the best method, reaching top-1 highest balanced accuracy and weighted accuracy in 8/12 scenarios. The PCA-kmeans baseline is second best with 4/12 top-1 ranking. RND underperforms by never achieving top-1 performance, yet it always ranks among the top-3 best methods. Additionally, most methods, including logpZO and RND, maintain practical detection times well below the average successful trajectory time. SPARC is on average the fastest as it attains top-1 performance in 4/6 cases, but it exhibits poor accuracy.
Overall, across all experiments, we find our novel learned logpZO score within the FAIL-Detect framework to be most consistent in performance. The post-hoc methods are often at the extremes of performance (either doing well or poorly) depending on the particular setting.
Qualitative score trends. Visualization of the detection scores (Figs. 6 and 9) confirms that learned methods are more discriminative with better score separation between successful and failed trajectories compared to post-hoc approaches. STAC also suffers from a single calibration threshold that is time invariant. In Section -D, we present comprehensive ablation studies examining performance sensitivity to CP significance level $\alpha$ (Fig. 10).